OpenStudy (shamil98):

using the trapezoidal rule. Just need to see where i went wrong.. \[\int\limits_{0}^{2} x^3 dx\] over 4 intervals so. n = 4 \[h = \frac{ b-a }{ n } = \frac{ 2-0 }{ 4 }= \frac{ 1 }{ 2 }\] \[T = \frac{ h }{ 2 } (y_0 + 2y_1 + 2y_2 +..... 2y_{n-1} + y_n)\] \[T = \frac{ \frac{ 1 }{ 2 } }{ 2 } (0 + 2(1) + 8) = \frac{ 1 }{ 4 } (10) = \frac{ 10 }{ 4 }\] Book answer is 17/4. So I did some mistake... but I don't see what I did wrong.

OpenStudy (anonymous):

Why do you only have \(3\) terms for \(y_0+y_1+y_2+y_3+y_4\)?

OpenStudy (shamil98):

oh, that's what I did wrong.

OpenStudy (shamil98):

\[T = \frac{ 1 }{ 4 }( 0 + 2(1) + 2(8) + 2(27) + 64)\] ?

OpenStudy (shamil98):

I thought it was only three terms, because it's x = 0 to x = 2. so 0 , 1 , 2 .

OpenStudy (campbell_st):

why not do it by repeated applications A = h/2(f(a) + f(a + h)) A = 1/2(f(0) + f(1/2)) + 1/2(f(1/2) + f(1)) + 1/2(f(1) + f(3/2)) + 1/2(f(3/2)+ f(2))

OpenStudy (shamil98):

I'm supposed to use the formula given from the book for this one.

OpenStudy (anonymous):

\[ y_k = y(kh+a) \]

OpenStudy (campbell_st):

well actually it is the formula is the book height/2 times sum of 1st + last + 2 times middle A = h/2[f(0) +2(f(1/2) + 2f(1) + 2f(3/2) + f(2)]

OpenStudy (shamil98):

Oh, I think I see what I did wrong. Since it's 4 sub intervals, so when x = 0, 0.5, 1,1.5, and 2. \[T = \frac{ 1 }{ 4 } ( 0 + 2(\frac{ 1 }{ 8 }) + 2(1) + 2(\frac{ 27 }{ 8 }) + 8)\] \[= \frac{ 1 }{ 4 }(\frac{ 2 }{ 8 } + 2 + \frac{ 54 }{ 8 } + 8)\] \[= \frac{ 1 }{ 4 }\left( 17 \right)\] \[ = 4.25\]